OpenStudy (anonymous):

simplify the expression if possible log 0.0001= The log is undefined.

OpenStudy (anonymous):

-4, son log 0.0001 = 10^x x=-4 or some such

OpenStudy (anonymous):

If the subscript on the LOG is undefined, you always assume it to be 10.

OpenStudy (anonymous):

those are two different options to the answer either it has an equal or its undefined

OpenStudy (unklerhaukus):

\[\log {0.0001}=\log10^{-4}=-4\log10=\]

OpenStudy (anonymous):

so putting this in the simpliest form would be -4?

OpenStudy (anonymous):

or log 10

OpenStudy (anonymous):

can I show you a screen shot?

OpenStudy (unklerhaukus):

one of those

OpenStudy (unklerhaukus):

use these \[\log a=\log_{10}a\] \[\log_bb=1\]

OpenStudy (anonymous):

that does not look like the ight format stand by please

OpenStudy (anonymous):

\[\log0.0001= \log _{10} 0.0001\]

OpenStudy (anonymous):

\[\log _{10}10^\left( -4 \right)\]

OpenStudy (anonymous):

-4 is the answer

OpenStudy (anonymous):

\[\log _{a^b}c^d=\frac{ d }{ b } \log _{a}c\]

OpenStudy (anonymous):

Thank you very much for the appreciation.